The prevailing market rates are as follows.

INR/USD = Rs. 77.00

Interest rate for a 6 month loan in India = 12% per annum

Interest rate for a 6 month loan in USA = 6% per annum

a. Explain the concept of Interest Rate Parity. What will be the expected 6 -months forward rate for US dollar in India?

b.Compute the Forward premium/discount of USD/INR in the Indian Forex Market?


1
Expert's answer
2021-11-03T10:45:53-0400

Solution:

a.). Interest rate parity (IRP) connects interest rates, spot exchange rates, and foreign exchange rates in foreign exchange markets.

The fundamental equation governing the relationship between interest rates and currency exchange rates is the IRP. The fundamental premise of IRP is that the hedged returns from investing in different currencies should be the same, regardless of their interest rates. IRP is the concept of no-arbitrage in the foreign exchange markets

 

Forward rate:

The formula rate formula (F0) = S0 x (1+ia/​1+ib​​)  

Where: F0​=Forward Rate

         S0 ​= Spot Rate = 77

         ia = Interest rate in country A (India)    = 12%                         

         ib​ = Interest rate in country B (United States) = 6%                                                                                                                                                                                                                                                                                                                                                             

Forward rate (F0) = 77 "\times" ("\frac{1 + 0.12}{1 + 0.06)}\times 100 = 77 \times \frac{1.12}{1.06} = 81.62"


Forward rate (F0) = 81.62


b.). Forward premium discount = "\frac{Forward\; rate-Spot\; rate}{Spot \; rate} \times 100"


FP = "\frac{81.62-77}{77} \times 100 = 6%"